# Low-energy $e^{+}\,e^{-}\to\gamma\,\gamma$ at NNLO in QED

### v1

@@ 1 Introduction
Diphoton production from a quark pair at NNLO in QCD has also received a lot of attention. Two-loop results with massive internal quarks and phenomenological results have been presented recently [14, 15, 16, 17, 18]. Even three-loop amplitudes with massless quarks have been considered [19]. There is, however, a crucial difference between QED and QCD calculations. In QED, the mass of the external fermions (i.e. the electrons) must not be neglected, whereas in all QCD computations, the external quarks are treated as massless. The difference originates from the fact that QED is more exclusive with respect to collinear emission.
@@ 3.1 Total cross sections
We have compared the NLO total cross section without any cuts to the analytic results given in [10]. For very low energies $\sqrt{s}\lesssim 10\,{\rm MeV}$ it is straightforward to obtain Monte Carlo errors of $10^{-6}$, whereas for energies approaching $\sqrt{s}\simeq 10\,{\rm GeV}$ the Monte Carlo errors increase to $10^{-4}$. In all cases we have found agreement within these errors.
@@
Our results are summarised in Table 1 and compared to [12]. The NLO results agree within 0.01% and the corrections are between 686-8%. Beyond NLO, there are physical differences between [12] and our results. We have included the full $\mathcal{O}(\alpha^{2})$ NNLO corrections, whereas in [12] photonic corrections beyond NLO have been implemented with a parton-shower approach. Thus, VP, LbL, and some non-logarithmic $\mathcal{O}(\alpha^{2})$ corrections are missing in [12]. On the other hand, they resum the leading-logarithmic corrections from multiple photon emission and, hence, include dominant terms of photonic corrections also beyond NNLO. Rather than comparing the best predictions it is thus more informative to compare the exact photonic NNLO corrections $\sigma^{(2,\gamma)}$ with the parton-shower induced contributions beyond NLO. In the notation of [12], where $\sigma^{\text{NLO}}_{\alpha}$ is used for $\sigma_{1}$, they are given by $\sigma_{\text{exp}}-\sigma_{1}$. As can be seen from Table 1, the two approaches agree within 102010-20% and give beyond-NLO photonic corrections of the order of 0.20.50.2-0.5%. Taking into account the $\mathcal{O}(\alpha^{2})$ suppression, this results in differences of the order of 0.05%.
@@ 3.2 Differential cross sections for a KLOE-like scenario
where $\theta_{l}$ and $\theta_{s}$ are the angles of the two most energetic photons satisfying (3). We apply a beam spread for both incoming particles, with a Gaussian distribution of FWHM 0.12 MeV [40]. A beam spread is able to cure largely fluctuating distributions caused by soft enhancements when the energies of tree-level photons are equal. Further, the sorting algorithm for photons in the final state requires a percent-level energy resolution, otherwise the event is randomly ordered.
@@ 3.3 Differential cross sections for a Belle-like scenario
Even though we use a symmetric angular cut, the parameters of the photons are understood to be in the LAB frame. Hence, photons tend to be emitted forward, i.e. $\theta_{\gamma}>90^{\circ}$. The centre-of-mass energy is $\sqrt{s}=10.583\,{\rm GeV}$. No beam spread is applied in this case, as the beam energies are asymmetrical by construction.

### v2

@@ 1 Introduction
In the context of processes involving photon pairs, it is worth mentioning available NLO electroweak calculations for $\gamma\gamma$ into light or heavy fermions [14, 15]. On the other hand, diphoton production from a quark pair at NNLO in QCD has also received a lot of attention. Following two-loop results with massless internal quarks [16, 17, 18, 19], calculations with massive internal quarks and phenomenological results have been presented recently [20, 21, 22, 23, 24]. Even three-loop amplitudes with massless quarks have been considered [25]. There is, however, a crucial difference between QED and QCD calculations. In QED, the mass of the external fermions (i.e. the electrons) must not be neglected, whereas in all QCD computations, the external quarks are treated as massless. The difference originates from the fact that QED is more exclusive with respect to collinear emission.
@@ 3 Results
The production runs employed version v0.7.0 of the McMule public release on Intel Xeon Gold 6140 CPUs. Timing details are specified for each configuration in the relevant sections. More details, e.g. the number of evaluation points, can be found in the repository of the McMule User Library
https://gitlab.com/mule-tools/user-library/-/tree/master/diphoton/
The major time bottleneck is due to NNLO photonic corrections, requiring 80–90% of the total computing time.
@@ 3.1 Total cross sections
We have compared the NLO total cross section without any cuts to the analytic results given in [10]. For very low energies $\sqrt{s}\lesssim 10\,{\rm MeV}$ it is straightforward to obtain Monte Carlo errors of $10^{-6}$, whereas for energies approaching $\sqrt{s}\simeq 10\,{\rm GeV}$ the Monte Carlo errors increase to $10^{-4}$. In all cases we have found agreement within these errors. In order to obtain the accuracy mentioned above, we ran McMule for 0.002 CPU-years for energies $\sqrt{s}\lesssim 100\,{\rm MeV}$, 0.3 CPU-years for energies $\sqrt{s}\gtrsim 1\,{\rm GeV}$.
@@
Our results are summarised in Table 1 and compared to [12]. Independently of the centre-of-mass energy, each configuration required a running time of 0.06 CPU-years. The NLO results agree within 0.01% and the corrections are between 686-8%. Beyond NLO, there are physical differences between [12] and our results. We have included the full $\mathcal{O}(\alpha^{2})$ NNLO corrections, whereas in [12] photonic corrections beyond NLO have been implemented with a parton-shower approach. Thus, VP, LbL, and some non-logarithmic $\mathcal{O}(\alpha^{2})$ corrections are missing in [12]. On the other hand, they resum the leading-logarithmic corrections from multiple photon emission and, hence, include dominant terms of photonic corrections also beyond NNLO. Rather than comparing the best predictions it is thus more informative to compare the exact photonic NNLO corrections $\sigma^{(2,\gamma)}$ with the parton-shower induced contributions beyond NLO. In the notation of [12], where $\sigma^{\text{NLO}}_{\alpha}$ is used for $\sigma_{1}$, they are given by $\sigma_{\text{exp}}-\sigma_{1}$. As can be seen from Table 1, the two approaches agree within 102010-20% and give beyond-NLO photonic corrections of the order of 0.20.50.2-0.5%. Taking into account the $\mathcal{O}(\alpha^{2})$ suppression, this results in differences of the order of 0.05%.
@@ 3.2 Differential cross sections for a KLOE-like scenario
where $\theta_{l}$ and $\theta_{s}$ are the angles of the two most energetic photons satisfying (3). We apply a beam spread for both incoming particles, with a Gaussian distribution of FWHM 0.12 MeV [46]. A beam spread is able to cure largely fluctuating distributions caused by soft enhancements when the energies of tree-level photons are equal. Further, the sorting algorithm for photons in the final state requires a percent-level energy resolution, otherwise the event is randomly ordered. The results obtained for the KLOE-like scenario required a running time of 0.07 CPU-years.
@@ 3.3 Differential cross sections for a Belle-like scenario
Even though we use a symmetric angular cut, the parameters of the photons are understood to be in the LAB frame. Hence, photons tend to be emitted forward, i.e. $\theta_{\gamma}>90^{\circ}$. The centre-of-mass energy is $\sqrt{s}=10.583\,{\rm GeV}$. No beam spread is applied in this case, as the beam energies are asymmetrical by construction. The results obtained for the Belle-like scenario required a running time of 0.7 CPU-years.